Diffusion Morning Report 08/01/07

Ted Berg tberg at stanford.edu
Wed Aug 1 06:58:11 PDT 2007


Good Morning All,
    Looking OK this morning. Ray is out on vacation for the next 2 
weeks. POCl3 is down for tube clean. TEOS2,BBr3 are under observation.  
LTO has a possible particle problem. FGA2 end cap stuck again.  No other 
issues at this time.

  	  	  	  	  	 
  	Date: 	8/1/07 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	yellow 	observation 	  	  	 
Tylan 6-POCl3 	down 	intermittant leak-tube clean 	  	  	 
Tylan 7-TEOS2 	yellow 	observation 	9.99/30 	  	 
Tylan 8-SiGe 	up 	  	19.49/30 	  	 
Tylan 9-Poly 	up 	  	16.97/30 	  	 
Tylan 10-Nitride 	up 	  	24.83/30 	  	 
Tylan 11-BPSG-LTO 	yellow 	particles? 	20.64/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	future 	  	
	
	
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	
	  	  	 
Burn Box 	Up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	yellow 	end cap stuck 	  	  	 
SiH4 for furnaces 	up 	A on line 640 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20070801/d4c12600/attachment.html>


More information about the diffusion mailing list